not in book. Factor each trinomial, if possible. If the trinomial cannot be factored using integers, white prime. I am checking my son's homework.
3a^2+30a+63
and
10w^2-19w-15

Answer by checkley75(3666)

3A^2+30A+63=0
(3A+9)(A+7)=0
3A+9=0
3A=-9
A=-9/3
A=-3 ANSWER.
A+7=0
A=-7 ANSWER.
------------------------------------------
10W^2-19W-15=0
(5W+3)(2W-5)=0
5W+3=0
5W=-3
W=-3/5 ANSWER.
2W-5=0
2W=5
W=5/2
W=2.5 ANSWER.